OpenStudy (anonymous):

Subtract and simplify 11 1/12 -2 2/3

OpenStudy (jdoe0001):

\(\bf 11\cfrac{1}{12}-2\cfrac{2}{3}\implies \cfrac{11\cdot 12+1}{12}-\cfrac{2\cdot 3+2}{3}=\cfrac{}{{\color{red}{ lcd?}}}\)

OpenStudy (anonymous):

8/3

OpenStudy (anonymous):

http://openstudy.com/study#/groups/History

OpenStudy (anonymous):

what is that